# Every nonflat conformal minimal surface is homotopic to a proper one

### v1

@@ 6.  Some remarks on minimal surfaces defined on compact
Assume that $D$ is a minimally convex domain in $\mathbb{R}^{3}$, and let $M$ be a compact bordered Riemann surface with nonempty boundary $bM$. Then, every nonflat conformal minimal immersion $u\colon M\to D$ is homotopic through nonflat conformal minimal immersions $\mathring{M}\to D$ to a proper complete nonflat conformal minimal immersion $\tilde{u}\colon\mathring{M}\to D$ with $\mathrm{Flux}(u)=\mathrm{Flux}(\tilde{u})$.
@@ 7.  Directed holomorphic curves
We point out that the choice of $1$-form $\theta$ is not important (as long as it is fixed) since the subvariety $A\subset\mathbb{C}^{n}$ is conical. A special case of directed $A$-immersions occurs when $A$ is the null quadric ${\bf A}_{*}$ (2.1); then a holomorphic $A$-immersion is precisely the null curve.
@@
Let $A\subset\mathbb{C}^{n}$ ($n\geq 3$) be a closed irreducible conical complex subvariety which is not contained in any hyperplane, is smooth away from 00, and such that $A_{*}=A\setminus\{0\}$ is an Oka manifold. Assume that for $k\in\{1,2\}$, the hyperplane section $A\cap\{z_{k}=1\}$ is an Oka manifold and the coordinate projection $\pi_{k}\colon A\to\mathbb{C}^{n}$ onto the $z_{k}$-axis admits a local holomorphic section $h_{k}$ near $z_{k}=0$ with $h_{k}(0)\neq 0$. Let $F\colon M\to\mathbb{C}^{n}$ be a nondegenerate holomorphic $A$-immersion from an open Riemann surface $M$ into $\mathbb{C}^{n}$. Then there exists a homotopy $F_{t}\colon M\to\mathbb{C}^{n}$, $t\in[0,1]$, of nondegenerate holomorphic $A$-immersions such that $F_{0}=F$ and $F_{1}$ is a proper nondegenerate holomorphic $A$-embedding.
@@
In analogy to Corollary 1.3 and Problem 1.4, we state the following consequence and an open problem.
@@
A positive answer to Problem 7.6 would, combined with [14, Theorem 5.6], tell us that the inclusion $\mathscr{I}_{A,*}^{p}(M,\mathbb{C}^{n})\lhook\joinrel\relbar\joinrel\rightarrow\mathscr{O}(M,{\bf A}_{*})$, or $\mathscr{I}_{A,*}^{c}(M,\mathbb{C}^{n})\lhook\joinrel\relbar\joinrel\rightarrow\mathscr{O}(M,{\bf A}_{*})$, respectively, is a weak homotopy equivalence. Equivalently, that would mean that the map $\mathscr{I}_{A,*}^{p}(M,\mathbb{C}^{n})\lhook\joinrel\relbar\joinrel\rightarrow\mathscr{C}(M,{\bf A}_{*})$, or $\mathscr{I}_{A,*}^{c}(M,\mathbb{C}^{n})\lhook\joinrel\relbar\joinrel\rightarrow\mathscr{C}(M,{\bf A}_{*})$, respectively, is a weak homotopy equivalence.

### v2

@@ 6. Some remarks on minimal surfaces defined on compact
Assume that $D$ is a minimally convex domain in $\mathbb{R}^{3}$, and let $M$ be a compact bordered Riemann surface with nonempty boundary $bM$. Then, every nonflat conformal minimal immersion $u\colon M\to D$ is homotopic through nonflat conformal minimal immersions $\mathring{M}\to D$ to a proper complete nonflat conformal minimal immersion $\tilde{u}\colon\mathring{M}\to D$ with $Flux(u~)\mathrm$$\mathrm{Flux}(\tilde{u})$.
@@ 7. Directed holomorphic curves
We point out that the choice of $1$-form $\theta$ is not important (as long as it is fixed) since the subvariety $A\subset\mathbb{C}^{n}$ is conical. A special case of directed $A$-immersions occurs when $A$ is the null quadric $\bf A$} (2.1); then a holomorphic $A$-immersion is precisely the null curve.
@@
Let $A\subset\mathbb{C}^{n}$ ($n\geq 3$) be a closed irreducible conical complex subvariety which is not contained in any hyperplane, is smooth away from 0, and such that $A_{*}=A\setminus\{0\}$ is an Oka manifold. Assume that for $k\in\{1,2\}$, the hyperplane section $A\cap\{z_{k}=1\}$ is an Oka manifold and the coordinate projection $\pi_{k}\colon A\to\mathbb{C}^{n}$ onto the $z_{k}$-axis admits a local holomorphic section $h_{k}$ near $z_{k}=0$ with $h_{k}(0)\neq 0$. Let $F\colon M\to\mathbb{C}^{n}$ be a nondegenerate holomorphic $A$-immersion from an open Riemann surface $M$ into $\mathbb{C}^{n}$. Then there exists a homotopy $F_{t}\colon M\to\mathbb{C}^{n}$, $t\in[0,1]$, of nondegenerate holomorphic $A$-immersions such that $F_{0}=F$ and $F_{1}$ is a proper nondegenerate holomorphic $A$-embedding.
@@
In analogy to Corollary 1.3 and Problem 1.5, we state the following consequence and an open problem.
@@
A positive answer to Problem 7.6 would, combined with [14, Theorem 5.6], tell us that the inclusion $\mathscr{I}_{A,*}^{p}(M,\mathbb{C}^{n})\lhook\joinrel\relbar\joinrel\rightarrow\mathscr{O}(M,A_{*})$, or $\mathscr{I}_{A,*}^{c}(M,\mathbb{C}^{n})\lhook\joinrel\relbar\joinrel\rightarrow\mathscr{O}(M,A_{*})$, respectively, is a weak homotopy equivalence. Equivalently, that would mean that the map $\mathscr{I}_{A,*}^{p}(M,\mathbb{C}^{n})\lhook\joinrel\relbar\joinrel\rightarrow\mathscr{C}(M,A_{*})$, or $\mathscr{I}_{A,*}^{c}(M,\mathbb{C}^{n})\lhook\joinrel\relbar\joinrel\rightarrow\mathscr{C}(M,A_{*})$, respectively, is a weak homotopy equivalence.